# Symplectic Neural Flows for Modeling and Discovery

### v1

@@ 4.2.1 Unsupervised experiments
Similarly to the harmonic oscillator example in Section 4.1, we solve the system’s equations of motion with the SympFlow architecture and with an unconstrained MLP. The MLP is again regularized by the additional loss term in Equation 11. In this section, we experiment training SympFlow with: i) Hamiltonian regularization Equation 12 and b) residual-based loss function Equation 9. We train both networks for $50,000$ epochs, setting $\Delta t=1$. Figure 6 show the damped harmonic oscillator solutions for both MLP and SympFlow, and for increasing values of the damping constant $\lambda$. We can see that both networks are able to capture the dissipative dynamics of the system for the different values of $\lambda$. This result is further supported by the left subfigure of Figure 9, where we also notice that adding an energy regularization term does not seem to improve SympFlow’s overall performance.
Figure 6: Damped harmonic oscillator solutions, where $q_{x}$ is the $1D$ position and $p_{x}$ is the (conservative) $1D$ momentum of the oscillator’s mass for different values of the damping constant $\lambda$, Appendix A. Top: $\lambda=0.1$, middle: $\lambda=0.5$, and bottom: $\lambda=1$.
@@ 4.2.2 Supervised experiments
In this section we compare the performance of the MLP network and SympFlow, trained with $N=50$ initial conditions and $M=10$ time samples for each of them. Figure 7, Figure 8, and Figure 9 show that SympFlow outperforms the MLP network when the dynamics of the system is more complex (case $\lambda=0.1$) and presents oscillations with decreasing amplitude, whereas for solutions with less complex behavior both methods seem to perform similarly well.
@@
Figure 8: SympFlow supervised experiment without regularization — Damped Harmonic Oscillator: Trained with $N=50$, $M=10$, and no noise, $\varepsilon=0$. procedure.
Figure 9: Comparison of the damped harmonic oscillator relative energy, normalised over the integration time, for unsupervised (left) and supervised (right) experiments. In the unsupervised case, both the MLP and SympFlow. In the supervised case, the solutions have been computed with $N=50$ initial conditions and $M=10$ sampling points. In this setting, with fewer initial conditions, SympFlow outperforms MLP.
@@ 4.3 Hénon–Heiles system
The Hénon–Heiles system is a model for studying non-linear dynamics, chaos, and the transition from regular to chaotic motion in physics. Initially developed to understand the motion of a star in a galactic potential, the system has since become a paradigmatic example in the study of chaos theory. The Hénon–Heiles system is a Hamiltonian system based on the Hamiltonian function [28, Section I.3]
@@
The equations of motion in Equation 25 provide a considerable challenge to numerical integrators and network-based simulations since they exhibit chaotic behavior corresponding to specific energy levels. Due to the chaotic nature of the system, it is unreasonable to expect a long-term agreement of the approximation with the exact solution. However, for chaotic systems, one would like to capture the correct global behavior of the trajectories. A common strategy to test this is to consider a two-dimensional Poincaré section and compare the obtained results for different methods. We proceed in this way to compare SympFlow with an MLP. To that end, we start by considering the Poincaré section corresponding to the linear subspace
@@
and we will provide a planar representation, in the variables $(q_{y},p_{y})$, of the intersection of the trajectories with the section $\mathcal{S}$. For a definition of the notion of attractor, and Poincaré sections, see [51].
We now move to the two experimental settings we consider for this system. All the experiments fix $\Delta t=1$ and $\Omega=[-1,1]^{4}\subset\mathbb{R}^{4}$. Some of these initial conditions correspond to energy levels leading to chaotic dynamics. However, since the training process only relies on the short integration interval $[0,1]$, this does not seem to negatively affect the quality of the recovered models.
@@ 4.3.1 Unsupervised experiments
As for the previous two test problems, we consider the unsupervised problem of solving the differential equations in Equation 25 for arbitrary initial conditions in $\Omega\subset\mathbb{R}^{4}$, and for time instants in $[0,\Delta t]$. The plots consider the initial condition $(0.3,-0.3,0.3,0)$, leading to initial energy $H_{0}=0.13$, corresponding to chaotic dynamics. We collect in Figure 10 the experiments with the three training regimes for SympFlow and with an MLP, again regularized by Equation 11. We see that SympFlow leads to better long-term energy behavior, and that the mixed training regime leads to slightly improved results. We also plot the Poincaré cuts associated with the section in Equation 26. A reference cut obtained with Runge–Kutta $(5,4)$ can be found in Appendix F. The SympFlows lead to much more qualitatively accurate solutions, given that the obtained cuts resemble the expected behavior. Appendix F also collects the plots of the solution curves obtained with this initial condition for the four trained models.
@@ 4.3.2 Supervised experiments
For the supervised experiment, we notice a pretty similar situation as in the unsupervised one, as can be seen in Figure 11. In this setup, it is also very hard to train the MLP model, as can be seen from the produced solution curves in Figure 13 of Appendix F. Once more, SympFlow demonstrates the ability to capture the correct global qualitative behavior of the solutions as we can see in both the energy plot and the Poincaré cuts in Figure 11.
Figure 11: Supervised experiment — Hénon–Heiles: Comparison of the MLP and the SympFlow trained with a dataset of parameters $N=100$, $M=50$, and $\varepsilon=0$. We show the energy plots for the time interval $[0,1000]$, and also the Poincaré sections obtained with the two different networks.
@@ 5 Conclusions and further work
In this work we presented SympFlow, a symplectic neural flow able to provide accurate and reliable long-term solutions for generic Hamiltonian systems. We have shown that SympFlow is a universal approximator and it can be applied to (i) approximately solve the equations of motion of Hamiltonian systems and (ii) approximate the flow map of unknown Hamiltonian system based on trajectory data. Because of its structure, SympFlow admits an underlying Hamiltonian which could also be used to study an approximated physical model, a task which has previously been studied in the scientific machine learning literature [3, 26]. Potentially, both SympFlow functionalities, (i)-(ii), can be combined to model complex physical systems. We have have also demonstrated the advantages of SympFlow over an unconstrained MLP in several experimental tasks, including a chaotic system and a non-conservative system.
The numerical experiments, show that the performance of the SympFlow is equal or superior to an MLP network for unsupervised tasks. For supervised tasks, however, SympFlow outperforms MLP to approximate complex solution, in particular when the number of initial conditions and time samples are significantly reduced, hence, highlighting the data efficiency of SympFlow.
There are several natural ways to extend this research. From the theoretical point of view, it would be interesting to further explore the energy behavior of the SympFlow since we experimentally observe a better error growth than the one in 2. We will also expand the applicability of SympFlow to higher-dimensional dynamical systems, such as spatially semi-discretized PDEs. To do so, the main effort will be in improving the computational efficiency of the SympFlow layers, which currently rely on automatic differentiation which can be costly for higher dimensional problems.
@@ Acknowledgements
The authors would like to thank Georg Maierhofer for useful discussions. DM acknowledges support from a grant from the Simons Foundation, the EPSRC Programme Grant on the Mathematics of Deep Learning (Maths4DL), and the Norwegian University of Science and Technology. CBS acknowledges support from the Philip Leverhulme Prize, UK, the Royal Society Wolfson Fellowship, UK, the EPSRC advanced career fellowship, UK EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards, UK 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No.777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information, UK and the Alan Turing Institute, UK. FS acknowledges support from the EPSRC advanced career fellowship EP/V029428/1. PC thanks the Alan Turing Institute for sponsoring her Daphne Jackson Fellowship. ZS acknowledges support from the Cantab Capital Institute for the Mathematics of Information and the Trinity Henry Barlow Scholarship scheme.

### v2

@@ 4.2.1 Unsupervised experiments
Similarly to the harmonic oscillator example in Section˜4.1, we solve the system’s equations of motion with the SympFlow architecture and with an unconstrained MLP. The MLP is not regularized in this situation, since the physical energy is not preserved for this system. In this section, we experiment training SympFlow with: i) Hamiltonian regularization Equation˜13 with $\gamma=1$ and b) residual-based loss function Equation˜10. We train both networks for $50,000$ epochs, setting $\Delta t=1$. Figure˜6 shows the damped harmonic oscillator solutions for both MLP and SympFlow, and for increasing values of the damping constant $\lambda$. We can see that both networks are able to capture the dissipative dynamics of the system for the different values of $\lambda$. This result is further supported by the left subfigure of Figure˜9, where we also notice that adding an energy regularization term does not seem to improve SympFlow’s overall performance.
Figure 6: Damped harmonic oscillator solutions, where $q_{x}$ is the $1D$ position and $p_{x}$ is the (conservative) $1D$ momentum of the oscillator’s mass for different values of the damping constant $\lambda$. Left: $\lambda=0.1$, center: $\lambda=0.5$, and right: $\lambda=1$.
@@ 4.2.2 Supervised experiments
In this section we compare the performance of the MLP network and SympFlow, trained with $N=50$ initial conditions and $M=10$ time samples for each of them. Figure˜7, Figure˜8, and Figure˜9 show that SympFlow outperforms the MLP network when the dynamics of the system is more complex (case $\lambda=0.1$) and presents oscillations with decreasing amplitude, whereas for solutions with less complex behavior both methods seem to perform similarly well.
@@
Figure 8: SympFlow supervised experiment — Damped Harmonic Oscillator: Trained with $N=50$, $M=10$, and no noise, $\varepsilon=0$.
Figure 9: Comparison of the damped harmonic oscillator relative energy, normalized over the integration time, for unsupervised (left) and supervised (right) experiments. In the unsupervised case, both the MLP and SympFlow. In the supervised case, the solutions have been computed with $N=50$ initial conditions and $M=10$ sampling points. In this setting, with fewer initial conditions, SympFlow outperforms MLP.
@@ 4.3 Hénon–Heiles system
The Hénon–Heiles system is a model for studying non-linear dynamics, chaos, and the transition from regular to chaotic motion in physics. Initially developed to understand the motion of a star in a galactic potential, the system has since become a paradigmatic example in the study of chaos theory. The Hénon–Heiles system is a Hamiltonian system based on the Hamiltonian function [29, Section I.3]
@@
The equations of motion in Equation˜28 provide a considerable challenge to numerical integrators and network-based simulations since they exhibit chaotic behavior corresponding to specific energy levels. Due to the chaotic nature of the system, it is unreasonable to expect a long-term agreement of the approximation with the exact solution. However, for chaotic systems, one would like to capture the correct global behavior of the trajectories. A common strategy to test this is to consider a two-dimensional Poincaré section and compare the obtained results for different methods. We proceed in this way to compare SympFlow with an MLP. To that end, we start by considering the Poincaré section defined as
@@
and we will provide a planar representation, in the variables $(q_{y},p_{y})$, of the intersection of the trajectories with the section $\mathcal{S}$. For a definition of the notion of Poincaré sections, see [57].
We now move to the two experimental settings we consider for this system. All the experiments fix $\Delta t=1$ and $\Omega=[-1,1]^{4}\subset\mathbb{R}^{4}$. Some of these initial conditions correspond to energy levels leading to chaotic dynamics. Since the training process relies on the short integration interval $[0,1]$, this does not seem to affect the quality of the recovered models negatively.
@@ 4.3.1 Unsupervised experiments
As for the previous two test problems, we consider the unsupervised problem of solving the differential equations in Equation˜28 for arbitrary initial conditions in $\Omega\subset\mathbb{R}^{4}$, and for time instants in $0,\Delta t]$. The plots consider the initial condition (0.3,0.3,0.3,0)(0.3,-0.3,0.3,0), leading to initial energy $H_{0}=0.13$, corresponding to chaotic dynamics. We collect in Figure˜10 the experiments with the two training regimes for SympFlow and an MLP, i.e., with and without regularization by Equation˜11 and Equation˜12 to promote Hamiltonian conservation. We see that SympFlow leads to better long-term energy behavior, and that the mixed training regime leads to slightly improved results, see Appendix˜G. We also plot the Poincaré cuts associated with the section in Equation˜29. A reference cut obtained with Runge–Kutta $(5,4)$ can be found in Appendix˜H. The SympFlows lead to much more qualitatively accurate solutions, given that the obtained cuts resemble the expected behavior. We also notice that the regularization term considerably improves the MLP predictions, but is not so significant for SympFlow, which captures the correct behavior without regularization. The regularized SympFlow performs slightly worse than the unregularized one, suggesting that it would be worth further exploring the design of a more suitable regularization term, or refining the balance between the two loss terms. Appendix˜H also collects the plots of the solution curves obtained with this initial condition for the four trained models.
@@ 4.3.2 Supervised experiments
For the supervised experiment, we notice a pretty similar situation as in the unsupervised one, as can be seen in Figure˜11. In this setup, it is also very hard to train the MLP model, as can be seen from the produced solution curves in Figure˜14 of Appendix˜H. Once more, SympFlow demonstrates the ability to capture the correct global qualitative behavior of the solutions as we can see in both the energy plot and the Poincaré cuts in Figure˜11.
Figure 11: Supervised experiment — Hénon–Heiles: Comparison of the MLP and the SympFlow trained with a dataset of parameters $N=100$, $M=50$, and $\varepsilon=0$. We show the energy plots for the time interval $[0,1000]$, and also the Poincaré sections.
@@ 5 Conclusions and further work
In this work we presented SympFlow, a symplectic neural flow able to provide accurate and reliable long-term solutions for generic Hamiltonian systems. We have shown that SympFlow is a universal approximator and it can be applied to (i) approximately solve the equations of motion of Hamiltonian systems and (ii) approximate the flow map of unknown Hamiltonian system based on trajectory data. Because of its structure, SympFlow admits an underlying Hamiltonian which could also be used to study an approximated physical model, a task which has previously been studied in the scientific machine learning literature [3, 27]. Potentially, both SympFlow functionalities, (i)-(ii), can be combined to model complex physical systems. We have have also demonstrated the advantages of SympFlow over an unconstrained MLP in several experimental tasks, including a chaotic system and a non-conservative system. The numerical experiments, show that the performance of the SympFlow is equal or superior to an MLP network for unsupervised tasks. For supervised tasks, however, SympFlow outperforms MLP, particularly when the number of initial conditions and time samples are significantly reduced, highlighting the data efficiency of SympFlow.
There are several natural ways to extend this research. From the theoretical point of view, it would be interesting to further explore the energy behavior of the SympFlow since we experimentally observe a better error growth than the one in ˜2. We will also expand the applicability of SympFlow to higher-dimensional dynamical systems, such as spatially semi-discretized PDEs. To do so, the main effort will be in improving the computational efficiency of the SympFlow layers, which currently rely on automatic differentiation which can be costly for higher dimensional problems. Additionally, since the Hamiltonian matching regularization term does not appear to benefit SympFlow, we will investigate how to better balance the two loss terms, as well as explore improved regularization strategies for training these constrained models.
@@ Acknowledgements
The authors would like to thank Georg Maierhofer for useful discussions, Lidia Gomes Da Silva for suggesting revisions to some of the plots, and the reviewers for their thoughtful comments on the universality proof, which made it much more rigorous. DM acknowledges support from a grant from the Simons Foundation, the EPSRC programme grant in ‘The Mathematics of Deep Learning’, under the project EP/V026259/1, and the Norwegian University of Science and Technology. CBS acknowledges support from the Philip Leverhulme Prize, UK, the Royal Society Wolfson Fellowship, UK, the EPSRC advanced career fellowship, UK EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards, UK 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No.777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information, UK and the Alan Turing Institute, UK. FS acknowledges support from the EPSRC advanced career fellowship EP/V029428/1. PC thanks the Alan Turing Institute for sponsoring her Daphne Jackson Fellowship. ZS acknowledges support from the Cantab Capital Institute for the Mathematics of Information and the Trinity Henry Barlow Scholarship scheme.
Appendix A Derivation of the long-time Hamiltonian system
Let us consider
and evaluate its derivative in time. We evaluate the following limit
$\dot{\psi}(t,x_{0})=\lim_{h\to 0}\frac{\psi(t+h,x_{0})-\psi(t,x_{0})}{h}$
$=\mathbf{J}\nabla\mathcal{H}(\bar{\psi})(t-c\Delta t,\psi(t,x_{0})).$
as written in Equation˜14.
Appendix B Proof of the a-posteriori result
Proof of ˜2.
To start, we notice that
$\mathcal{H}(\psi)(t,x)$ $=\mathcal{H}(\bar{\psi})(t-\Delta t\lfloor t/\Delta t\rfloor,x)=H(x)+\delta(t-\Delta t\lfloor t/\Delta t\rfloor,x)$
and since $\mathbf{J}$ preserves the $\ell^{2}$ norm, we get that $\|\nabla\delta(t,x)\|_{2}\leq\varepsilon_{2}$ too. This implies that
where $C$ upper bounds the Lipschitz constant of $H$ over $\Omega$. Thus, $\left|H(\psi_{t}(x))-H(x)\right|\leq c_{2}\varepsilon_{2}t$. Moving to the variation of $\mathcal{H}(\psi)$, we notice that
$\mathcal{H}(\psi)(t,\psi_{t}(x))-\mathcal{H}(\psi)(0,x)$ $=\mathcal{H}(\psi)(t,\psi_{t}(x))-H(\psi_{t}(x))+H(\psi_{t}(x))-H(x)$
$+H(x)-\mathcal{H}(\psi)(0,x),$
which allows us to conclude
Appendix C Description of the network architectures
In what follows, we provide details on the network architectures considered in Section˜4. All the networks we compare have the same number of layers, where each layer is defined as a transformation acting on both parts, $q$ and $p$, which partition the phase-space variable $x\in\mathbb{R}^{2d}$. The number of layers chosen in the numerical experiments in Section˜4 is based on SympFlow’s performance. Thus, in the unsupervised experiments, the number of layers is set to three, whereas in the supervised experiments is set to five. Additionally, in all experiments the activation function is $\sigma=\tanh$.
C.1 SympFlow
Given a SympFlow, the $i$-th layer of $\bar{\psi}_{t}$ is given by the composition map $\phi_{\mathrm{p}}^{i}\circ\phi_{\mathrm{q}}^{i}$. Each of these two composed maps depends on network weights only via a potential function, which is denoted as $V_{\mathrm{q}}^{i}$ and $V_{\mathrm{p}}^{i}$. We model both potential functions with a feedforward neural network. We explicit it for $V_{\mathrm{q}}^{i}$, and the same applies to $V_{\mathrm{p}}^{i}$, but with different weights:
[…]
where $\ell_{1}:\mathbb{R}^{d+1}\to\mathbb{R}^{h}$, $\ell_{2}:\mathbb{R}^{h}\to\mathbb{R}^{h}$, $\ell_{3}:\mathbb{R}^{h}\to\mathbb{R}$, $h=10$, are three parametrized affine maps of the form $\ell_{j}(x)=A_{j}x+b_{j}$ for suitably shaped weight $A_{j}$ and bias $b_{j}$, $j=1,2,3$. For the case of the simple harmonic oscillator, where $d=1$, we thus have $30$ parameters modeling $\ell_{1}$, $110$ modeling $\ell_{2}$, and $11$ modeling $\ell_{3}$. A SympFlow of $5$ layers will hence have $5\cdot 2\cdot(30+110+11)=1510$ parameters.
C.2 MLP
for an affine map parametrized as $\ell_{i}(z)=A_{i}z+b_{i}$, $A_{i}\in\mathbb{R}^{c_{i+1}\times c_{i}}$ and $b_{i}\in\mathbb{R}^{c_{i+1}}$. For the first layer we have $c_{i}=2d+1$, since $z=[x,t]$, whereas for the last layer $c_{L+1}=2d$. All the intermediate dimensions are fixed to $c_{2}=...=c_{L}=10$. For the simple harmonic oscillator, considering an MLP of $5$ layers, we thus have that the first layer is parametrized by $40$ parameters, the intermediate three by $110$ each, and the last by $22$. This leads to a network with $40+330+22=392$ parameters.
Since composing maps as in Equation˜30 does not allow to enforce the initial condition, we modify the composition and define the MLP network as
[…]
Appendix D Background on symplectic maps and Hamiltonian dynamics
This section provides some background material on symplectic maps, Hamiltonian systems, and symplectic numerical methods.
Definition 1.
A differentiable map $F:\Omega\to\mathbb{R}^{2d}$, $\Omega\subseteq\mathbb{R}^{2d}$ open, is symplectic if, for every $x\in\Omega$, $(F^{\prime}(x))^{\top}\mathbf{J}F^{\prime}(x)=\mathbf{J}$.
Lemma 1.
Let $F,G:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ be two differentiable symplectic maps. Then, the maps $F\circ G,G\circ F:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ are symplectic as well.
Proof.
We prove the result for $F\circ G$, and the argument repeats analogously for $G\circ F$. By the chain rule, it holds that
$(F\circ G)^{\prime}(x)=F^{\prime}(G(x))G^{\prime}(x).$
Thus,
$((F\circ G)^{\prime}(x))^{\top}\mathbf{J}(F\circ G)^{\prime}(x)=(G^{\prime}(x))^{\top}(F^{\prime}(G(x)))^{\top}\mathbf{J}F^{\prime}(G(x))G^{\prime}(x)=(G^{\prime}(x))^{\top}\mathbf{J}G^{\prime}(x)=\mathbf{J}$
as desired. ∎
We then prove that the time-$t$ flow of a time-dependent Hamiltonian system is a symplectic diffeomorphism if the Hamiltonian function is twice continuously differentiable.
We recall that a time-dependent Hamiltonian system is described by the system of differential equations
$\dot{x}(t)=\mathbf{J}\nabla_{x}H(t,x(t)),$
for a time-dependent Hamiltonian energy function $H:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$. Such a definition could also be restricted to an open subset $\Omega\subseteq\mathbb{R}^{2d}$. We remark that the notation $\nabla_{x}$ stands for the gradient with respect to the second entry.
Proposition 2.
Let $H:\mathbb{R}\times\Omega\to\mathbb{R}$ be a twice continuously differentiable function on $\Omega\subseteq\mathbb{R}^{2d}$ open. Then, for each fixed $t\in\mathbb{R}$, the timet-t flow map $\phi_{H,t}:\Omega\to\mathbb{R}^{2d}$ is symplectic whenever defined.
Proof.
We recall that the flow map $\phi_{H,t}:\Omega\to\mathbb{R}^{2d}$ satisfies
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{H,t}(x_{0})=\mathbf{J}\nabla_{x}H\left(t,\phi_{H,t}(x_{0})\right)$
for every $t\geq 0$ and $x_{0}\in\Omega$. Differentiating both sides with respect to $x_{0}$, we get
$\frac{\partial}{\partial x_{0}}\frac{\mathrm{d}}{\mathrm{d}t}\phi_{H,t}(x_{0})=\mathbf{J}\nabla^{2}_{x}H(t,\phi_{H,t}(x_{0}))\,\frac{\partial}{\partial{x_{0}}}\phi_{H,t}(x_{0}),$
where $\nabla^{2}_{x}H$ is the Hessian matrix of $H$ with respect to the second component. Changing the differentiation order on the left, and calling $S_{x_{0}}(t)=\partial_{x_{0}}\phi_{H,t}(x_{0})$, we see that
$\frac{\mathrm{d}}{\mathrm{d}t}S_{x_{0}}(t)=\mathbf{J}\nabla^{2}_{x}H(t,\phi_{H,t}(x_{0}))S_{x_{0}}(t).$ (31)
We can then compute
$\frac{\mathrm{d}}{\mathrm{d}t}\left(S_{x_{0}}(t)^{\top}\mathbf{J}S_{x_{0}}(t)\right)=\left(\frac{\mathrm{d}}{\mathrm{d}t}S_{x_{0}}(t)\right)^{\top}\mathbf{J}S_{x_{0}}(t)+S_{x_{0}}(t)^{\top}\mathbf{J}\left(\frac{\mathrm{d}}{\mathrm{d}t}S_{x_{0}}(t)\right)$
$=\left(S_{x_{0}}(t)^{\top}\nabla^{2}_{x}H(t,\phi_{H,t}(x_{0}))\mathbf{J}^{\top}\right)\mathbf{J}S_{x_{0}}(t)+S_{x_{0}}(t)^{\top}\mathbf{J}\left(\mathbf{J}\nabla^{2}_{x}H(t,\phi_{H,t}(x_{0}))S_{x_{0}}(t)\right).$
Since $\mathbf{J}^{\top}\mathbf{J}=I_{2d}=-\mathbf{J}^{2}$ we conclude that the quantity above vanishes and hence
$S_{x_{0}}(t)^{\top}\mathbf{J}S_{x_{0}}(t)=S_{x_{0}}(0)^{\top}\mathbf{J}S_{x_{0}}(0).$
At time $t=0$, we recall that
$S_{x_{0}}(0)=\partial_{x_{0}}\phi_{H,0}(x_{0})=\partial_{x_{0}}x_{0}=I_{2d}$
which allows us to conclude $S_{x_{0}}(t)^{\top}\mathbf{J}S_{x_{0}}(t)=\mathbf{J}$ for every $t\geq 0$ as desired. ∎
To numerically approximate the solutions of a dynamical system, one needs to use numerical methods. When the target system possesses specific qualitative properties, such as energy preservation, symplecticity, or volume preservation, it is desirable to rely on numerical methods that reproduce these properties. Designing methods compatible with the qualitative properties of dynamical systems is the primary goal of an area of numerical analysis known as geometric numerical integration, as described in [29]. Symplectic one-step numerical methods are defined by maps $\varphi^{t_{0},\Delta t}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$, $\varphi^{t_{0},\Delta t}(x(t_{0}))\approx x(t_{0}+\Delta t)$, which, when applied to a Hamiltonian system, are symplectic maps. The Lie-Trotter splitting method is an example of a symplectic method, also often referred to as the symplectic Euler method in the context of symplectic integration. We utilise this scheme in our proof of the universal approximation theorem, so we briefly introduce it here. Consider a time-dependent separable Hamiltonian function, i.e., a function $H:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ of the form $H(t,(q,p))=K(t,p)+V(t,q)$ where $K,V:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$ could be thought as the kinetic and potential energies of the system. Here, we assume $x=(q,p)\in\mathbb{R}^{2d}$ is split into two components of the same dimension $q,p\in\mathbb{R}^{d}$. We can then write the Hamiltonian vector field of $H$ as the sum of two simpler Hamiltonian systems:
$\dot{x}(t)=\mathbf{J}\nabla_{x}H(t,x(t))=\begin{bmatrix}\nabla_{p}K(t,p(t))\\ 0_{d}\end{bmatrix}+\begin{bmatrix}0_{d}\\ -\nabla_{q}V(t,q(t))\end{bmatrix}$.
The two components on the right-hand side are Hamiltonian systems that can be exactly solved up to being able to integrate the gradients of the two components $K$ and $V$:
$\phi_{K,t_{0},\Delta t}(q_{0},p_{0})$ $=\left(q_{0}+\int_{t_{0}}^{t_{0}+\Delta t}\nabla_{p}K(t,p_{0})dt,\;p_{0}\right),$
$\phi_{V,t_{0},\Delta t}(q_{0},p_{0})$ $=\left(q_{0},\;p_{0}-\int_{t_{0}}^{t_{0}+\Delta t}\nabla_{q}V(t,q_{0})dt\right).$
The notation $\phi_{K,t_{0},\Delta t}(q_{0},p_{0})$ and $\phi_{V,t_{0},\Delta t}(q_{0},p_{0})$ is used to express the solution at time $t_{0}+\Delta t$ of the two Hamiltonian initial value problems with Hamiltonians $K$ and $V$, respectively, starting from time $t_{0}$. The two maps in the previous equation are exactly the layers of our SympFlow. The symplectic Euler method, applied to approximate $\phi_{H,t_{0},\Delta t}$ is thus defined as
$\varphi^{t_{0},\Delta t}=\phi_{K,t_{0},\Delta t}\circ\phi_{V,t_{0},\Delta t},$
and it could also be written by composing the flows in the opposite order, despite generating a different map, which still goes under the same name. We note that if $H$ is time-independent, then the integrals are easily solvable, and the flows we are composing reduce to linear shifts in one of the two components.